OpenStudy (mackenzie2013):

There are 6.02 × 1023 atoms in 1 gram of hydrogen. How many atoms are there in 800 grams of hydrogen? Write your answer in scientific notation. If necessary, round your answer to the nearest hundredth.

OpenStudy (campbell_st):

well 800 can be written as \[8 \times 10^2\] so the problem is now \[(6.02 \times 10^{23}) \times 800~~~~or~~~(6.02 \times 10^{23}) \times 8 \times 10^2 = \]